OpenStudy (superflytyguy):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12

OpenStudy (superflytyguy):

Here are the choices. A. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 B. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 C. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 D. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9

OpenStudy (superflytyguy):

I am so stuck on this

OpenStudy (uri):

y=mx+b formula ^ so arrange the equation first

OpenStudy (uri):

-11x + 9y = -12 9y=11x-12

OpenStudy (uri):

y=11/9x-12/9 y=11/9x-4/3

OpenStudy (uri):

@SuperflyTyGuy Can you tell now?

OpenStudy (superflytyguy):

Yea I think so, you translated it to y =mx + b form then divided 11 on both sides, but how did you get 12 to divide by 9? I see you simplified it to get 4/3

OpenStudy (superflytyguy):

@uri

OpenStudy (uri):

It was 9y=11x-12 right?

OpenStudy (superflytyguy):

right

OpenStudy (uri):

shifted 9 and divided it with 11x and -12

OpenStudy (uri):

y=11/9x-12/9 get it?

OpenStudy (superflytyguy):

ok i got it now, sorry I'm so tired i'm not thinking straight.

OpenStudy (uri):

So what's the answer?

OpenStudy (superflytyguy):

y = 11/9x - 4/3

OpenStudy (superflytyguy):

B

OpenStudy (uri):

A

OpenStudy (uri):

A. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3

OpenStudy (uri):

B has y intercept as 4/3 not -4/3.

OpenStudy (superflytyguy):

I always miss the negative signs in math. You're right y-intercept: -4/3

OpenStudy (superflytyguy):

4/3

OpenStudy (uri):

-4/3 not 4/3 god lol

OpenStudy (superflytyguy):

Thank you @uri

OpenStudy (uri):

Is this Joel?

OpenStudy (superflytyguy):

No. It's Ty, as in SuperflyTyGuy

OpenStudy (uri):

Alright. :D

OpenStudy (superflytyguy):

Thank you so much! It's really early and I'm so tired, I can't think straight

OpenStudy (uri):

yw. :)